OpenStudy (anonymous):

Use the limit definition to compute f′(a) and find an equation of the tangent line. f(t) = t − 10 t^2, a = 3. Y=

OpenStudy (tkhunny):

What limit definition would be to your liking?

OpenStudy (anonymous):

Umm the easiest one! (:

OpenStudy (anonymous):

doesnt it have to be f(x+x)-f(x)/x

OpenStudy (tkhunny):

I'm sure you mean "f(x+h)" and "/h". I'm also sure you mean \(\dfrac{f(x+h) - f(x)}{h}\), but this is not what you have written. Given the first two corrections, you have written \(f(x+h) - \dfrac{f(x)}{h}\), and this is absolutely not what you intend. Please remember your Order of Operations! Okay, having said that, you'll need: f(t) - which you have. f(t+h) - Please show this in it's expanded form.

OpenStudy (tkhunny):

Okay, make that f(3) and f(3+h).

OpenStudy (anonymous):

Yea I meant the top equation, what do you mean by expanded form?

OpenStudy (anonymous):

f(3):(3-10(3)^2

OpenStudy (tkhunny):

f(3) = 3 - 10(3)^2 = 3 - 10(9) = 3 - 90 = -87 f(3+h) = (3+h) - 10(3+h)^2 = 3 + h - 10(9 - 6h + h^2) = 3 + h - 90 + 60h - 10h^2 f(3+h) = -87 + 61h - 10h^2 Okay, now assemble the numerator and simplify.

OpenStudy (anonymous):

10(3)^2-61(3)-87=6

OpenStudy (anonymous):

or do i do the derivative and then plug in?

OpenStudy (tkhunny):

I don't know what that is. f(3+h) - f(3) = ( -87 + 61h - 10h^2) - (-87) Simplify.

OpenStudy (anonymous):

f(3+h)-f(3)= 7569+5407h-870h^2 or is it 61h-10h^2?

OpenStudy (tkhunny):

I am not encouraged by that. You should know the difference between multiplication and subtraction. It is the latter. Okay, now we have the numerator, let's go get the denominator and simplify again.

OpenStudy (anonymous):

Sorry my computer would not load, am still confused on what to do.

OpenStudy (anonymous):

should be \[f'(a)=\lim_{x \rightarrow a}\frac{ f(x-a)-f(a) }{ x-a }\]

OpenStudy (anonymous):

would it be, (t-10t^2-3)-(3)/(t − 10 t^2-3)?

OpenStudy (anonymous):

\[\lim_{t \rightarrow 3}\frac{ f(t-3)-f(3) }{ t-3 }\]

OpenStudy (anonymous):

what's f(t-3) = ? what's f(3) = ? figure these out and plug them in. then you'll need to factor out (t-3)

OpenStudy (anonymous):

What do I plug in for t?

OpenStudy (anonymous):

in the first one?

OpenStudy (anonymous):

\[f(t-3)=(t-3)-10(t-3)^2\]

OpenStudy (anonymous):

I got an answer of 3

OpenStudy (anonymous):

or do you want -10 t^3+90 t^2-270 t+270

OpenStudy (anonymous):

sorry it's like this... \[f'(a)=\lim_{x \rightarrow a}\frac{ f(x)-f(a)}{ x-a }\] so we get...\[f'(3)=\lim_{x \rightarrow 3}\frac{ (x-10x^2)-(3-10(3)^2) }{x-3 }=\lim_{x \rightarrow 3}\frac{ (x-10x^2)+87 }{x-3 }\] \[=\lim_{x \rightarrow 3}\frac{ (x-3)(-10x-29) }{x-3 }==\lim_{x \rightarrow 3}-10x-29=-59\]

OpenStudy (anonymous):

ok I see what you did, however when i type in -59 as the answer in the website it comes up at incorrect?

OpenStudy (anonymous):

hold on, perhaps I miscalculated... nope that's it. f(t) = t-10t^2, f'(t) = 1-20t... f'(3) = 1 - 20(3) = 1 - 60 = -59

OpenStudy (anonymous):

that's just the slope of the tangent line... it's not the equation. now that you have the slope, it has to go through the point (3, -87). use these facts to write the equation of the line with slope = -59, going through the point (3, -87)

OpenStudy (anonymous):

oh thats it I have to use an equation y=-59x-87 right

OpenStudy (anonymous):

no... when x = 3 y = -87... that's not the y intercept. this is calculus. you should have linear equations mastered!

OpenStudy (anonymous):

lol hold on let me try this again

OpenStudy (anonymous):

y+87=-59(x-3)

OpenStudy (anonymous):

y=-146x+177

OpenStudy (anonymous):

grrr now it says this Entered Answer Preview Result Messages -146x+177 incorrect Variable 'x' is not defined in this context

OpenStudy (anonymous):

the slope is wrong anyways... it's not -146, it's -59. as for how the system takes it, I can't help you there.

OpenStudy (anonymous):

o i finally get why it was wrong instead of x its t, thanks for all the help!